# Efficient Deep Demosaicing with Spatially Downsampled Isotropic Networks

### v1

@@ 1 Introduction
Demosaicing is a first and crucial step in an Image Sensor Processing (ISP) pipeline [Karaimer]. Recent advances in mobile photography, including more challenging non-Bayer CFA mosaics and pixel binning techniques, have made the task of demosaicing in modern mobile phones increasingly difficult [ESUM]. Thus, deep learning techniques for non-Bayer demosaicing have been increasingly explored [BJDD, SAGAN, ESUM]. However, deep learning techniques are infamously compute-heavy, which poses challenges for usage on a mobile platform. Consequently, lightweight networks are essential for demosaicing on mobile phones [BMTNet].
One major class of image demosaicing network are isotropic networks, which are often called residual-in-residual networks [JDNDMSR, ESUM, SGNet]. These networks are characterized by their entire trunk maintaining the same spatial resolution; one particularly well-known example is the Vision Transformer (ViT) architecture[ViT]. However, unlike ViT networks, which perform aggressive patch-ification to reduce computational cost, isotropic networks for demosaicing largely perform their computations on the full spatial resolution of the image (or nearly so), consequently leading to a comparatively higher computational cost. This particular property of isotropic image demosaicing networks seems to be inherited from the strong success of isotropic networks in super-resolution[RCAN, SwinIR, X-Restormer]. On the contrary, for the task of image demosaicing, we find that spatial downsampling is effective for improving performance and efficiency in FLOP-equivalent network variants.
Reasoning about the effect of singular variables in designing deep networks can often be challenging, due to the dozens of interdependent hyperparameters found in network architectures. We employ two measures to enable empirical testing of the effects of downsampling. First, we consider only very simple fully-convolutional isotropic networks, which reduces our search space to only three variables: depth, width, and downsampling factor. Next, to produce principled and fair candidate variants of downsampled and non-downsampled networks, we employ a mathematical architecture design technique modified from DeepMAD[DeepMAD]. Our results show that downsampling improves the performance of two FLOP-equivalent networks, especially in particularly small networks.
@@
Recent work in image demosaicing has explored complex patterns[BJDD, SAGAN, BMTNet] and techniques to create unified models[KLAP, ESUM]. Following this diversity in recent image demosaicing work, we test the applicability of downsampled networks to a variety of tasks. Our downsampled networks, which we term JD3Net (Joint Demosaicing and Denoising with Downsampled Network), are highly simple fully convolutional networks, but nonetheless we find that they achieve state-of-the-art efficiency on unified Bayer/non-Bayer joint-demosaicing-and-denoising, classical Bayer image demosaicing, and quad-Bayer HybridEVS image demosaicing (see Fig. 1). An overview of the CFA patterns considered can be found in Fig. 2. Overall, we argue that downsampling can broadly improve the efficiency of deep networks across a variety of image demosaicing tasks.
@@ 2.1 Tasks in Demosaicing
Joint demosaicing and denoising (JDD) is a deep learning task that combines two image processing steps – demosaicing and denoising – into one end-to-end objective. JDD is attractive because it preserves the decorrelated noise of raw images while maintaining the full information of the raw image for demosaicing[DeepJoint]; consequently, research in this direction has existed even before the popularization of deep learning [Klatzer]. Nonetheless, deep learning approaches have been the mainstream in JDD research recently. Typically, a convolutional neural network (CNN) variant is applied to this task [SGNet, BJDD, SAGAN, ESUM, JDNDMSR], with some exceptions [DemosaicFormer, BMTNet].
While early JDD research was largely based on the traditional Bayer CFA [JDNDMSR, SGNet, TENet], modern research often focuses on non-Bayer demosaicing[BJDD, SAGAN, TSANet, BMTNet] or unified Bayer/non-Bayer demosaicing[KLAP, ESUM]. Non-Bayer CFAs tend to be more challenging, and can even have pixels with completely zero information, such as in HybridEVS image demosaicing. The difficulty of non-Bayer demosaicing, even without the denoising task, similarly merits deep learning.
@@ 2.2 Isotropic Networks for Demosaicing
The neural networks applied to demosaicing can mainly be subdivided into two types of macro-architectures: U-Net [BJDD, SAGAN, DemosaicFormer, BMTNet, KLAP, TSANet] and Isotropic (often called residual-in-residual)[SGNet, JDNDMSR, ESUM, DeepJoint, RNAN, GRL].
In general, the main difference between U-Net and isotropic macro-architectures lies in downsampling. U-Net networks employ progressive downsampling to capture multiple feature-scales [UNet], while the trunks of isotropic networks employ a uniform spatial size [ESUM, JDNDMSR, RCAN], and employ spatial downsampling or upsampling only at the beginning or end of the network.
For most general image restoration tasks, such as image denoising or super-resolution, an isotropic network employs no downsampling at all [RCAN, SwinIR]. For demosaicing, while many isotropic networks also do not perform downsampling [ESUM, GRL, RNAN], some networks perform ”packing”, which folds a CFA pattern into a single pixel[DeepJoint, JDNDMSR]. However, prior networks do not explicitly use downsampling to improve efficiency and do not exceed 2x downsampling. A brief summary of prior isotropic networks for demosaicing is shown in Tab. 1.
@@
DeepJoint[DeepJoint] 2016 2 14
MMNet[MMNet] 2018 1 18
RNAN [RNAN] 2019 1 162
TENet [TENet] 2019 2 3199
SGNet [SGNet] 2020 2 -
JDNDM[JDNDMSR] 2021 2 408
GRL[GRL] 2023 1 491
ESUM[ESUM] 2025 1 408
@@
Because isotropic networks tend to operate on high spatial resolutions, without employing spatial downsampling, they are often quite slow and inefficient[X-Restormer]. This is a notable problem for demosaicing, since its main application field is mobile phones, which are famously compute constrained. In particular, isotropic networks for JDD can be several hundred G-FLOPs for 256-by-256 pixel images[ESUM, JDNDMSR], which are several orders of magnitude larger than networks typically designed for image recognition on mobile devices[MBv2, MBv3].
@@ 2.3 Zero-Shot Neural Architecture Search
The primary tool we will use to make this argument is Zero-Shot Neural Architecture Search (NAS). In contrast to conventional NAS techniques, Zero-Shot NAS attempts to find a performant neural network architecture without any training, often based on mathematical observations regarding the nature of deep neural networks[DeepMAD, Zen-NAS, MAE-Det]. This is attractive for designing networks in a principled manner, and is the primary reason we employ Zero-Shot NAS techniques in this paper.
Often, Zero-Shot NAS techniques employ the Principle of Maximum Entropy, which aims to maximize the entropy of the feature map outputted by the network [DeepMAD, Zen-NAS]. In particular, this paper focuses on the DeepMAD[DeepMAD] entropy formulation, which is designed for convolutional neural networks and takes the following form:
@@ 3.1 Network Design
JD3Net employs modified variants of NAFBlocks[NAFNet], with the attention removed, and which we will refer to as Simplified-NAFBlocks. A specific illustration of NAFBlocks and Simplified-NAFBlocks can be found in Fig. 3.
Our decision to remove channel attention from JD3Net is due to the implementation complexity of attention in image restoration networks. For instance, techniques such as Test-time Local Converter[TLC] are used by some networks, but not others. Moreover, we provide some brief ablation in Sec. 4.5, which shows that the Simple Channel Attention (SCA) mechanism used by NAFNet actually decreases the performance of JD3Net. Since JD3Net can acheive strong empirical performance without complex attention mechanisms, we leave the investigation of attention for downsampled isotropic networks to future work.
@@ 3.2 Mathematical Architecture Design
As previously stated, we modify DeepMAD’s entropy score to design JD3Net. The principal issue with applying DeepMAD directly to image restoraiton is that its entropy score is reliant on the spatial resolution of the input image (ie a network would have a different entropy score with a 256x256 image compared to a 512x512 image). Consequently, DeepMAD selects different networks depending on the input resolution. For classification, where input resolution is an active consideration in network design [EfficientNet, ConvNext], this makes sense. But for image restoration, where networks should be able to process an arbitrary image, we would prefer our selection of network to be invariant to the input resolution.
@@ 4.1 Searching for JD3Net
We perform our search for JD3Net by comparing validation PSNRs on the Hard Demosaicing Dataset[ESUM], a dataset of 638 real world joint-demosaicing-and-denoising images from 17 difficult scenes. Specifically, we compare results on the most noisy setting, ISO 3200. We train for 150 epochs using the provided pre-processed 48x48 hard training patches. Learning rate is $0.001$ on Adam optimizer with MSE loss (following ESUM[ESUM]). Batch size for the smaller network is 64, and 48 for the larger network.
@@
JDNDM[JDNDMSR] 408 53.69 - - 52.34 - - 50.43 - - 49.00 - - -
BJDD[BJDD] 69 - 50.86 - - 50.05 - - 48.88 - - 47.50 - -
SAGAN[SAGAN] 342 - - 49.55 - - 49.06 - - 48.05 - - 46.88 -
KLAP[KLAP] 106 53.27 52.01 50.44 51.91 50.99 49.79 50.28 49.40 48.59 48.95 48.10 47.20 50.08
ESUM[ESUM] 408 53.75 52.68 51.96 52.17 51.36 50.76 50.64 50.01 49.46 48.98 48.57 48.11 50.70
@@
BMTNet*[BMTNet] 7 37.69/0.980 34.79/0.950 36.11/0.981 34.45/0.973 33.95/0.965 35.398/0.970
NAFNet[NAFNet](w-32) 32 38.60/0.980 36.18/0.961 37.12/0.985 35.63/0.978 35.24/0.972 36.554/0.975
TSANet-S[TSANet] 37 38.73/0.984 - 36.56/0.984 36.15/0.980 35.19/0.973 36.658/0.980
@@
Restormer[Restormer] 282 39.16/0.986 36.54/0.967 37.11/0.985 36.36/0.977 35.00/0.971 36.83/0.977
SAGAN[SAGAN] 342 36.14/0.959 32.58/0.939 30.53/0.931 29.89/0.946 28.22/0.917 31.47/0.938
TSANet-L[TSANet] 149 39.40/0.986 - 37.34/0.986 37.07/0.983 35.76/0.977 37.393/0.983
DemosaicFormer[DemosaicFormer] 491 39.32/0.982 37.88/0.963 37.65/0.982 37.64/0.980 34.86/0.986 37.47/0.979
@@ 4.2 Real Image Unified Joint-Demosaicing-and-Denoising
Averaged across all CFA patterns and all noise levels, JD3Net-S achieves only 0.1 PSNR less than ESUM[ESUM] while being 16.3 times faster, while JD3Net outperforms ESUM by 0.35 PSNR while being 3.2 times faster. We show a selection of qualitative results between JD3Net and ESUM in Fig. 4, which demonstrates the better reconstruction capability of JD3Net. Quantitative results are given in Tab. 4.
@@ 4.3 Synthetic Quad-Bayer HybridEVS Demosaicing
We also evaluate JD3Net on synthetic quad-bayer HybridEVS demosaicing. We train on the MIPI dataset [MIPI] training split for $8\cdot 10^{6}$ iterations at a batch size of 20. We use PSNR loss[NAFNet] and Adam optimizer with a learning rate of $0.001$. We apply the data augmentations used by DemosaicFormer[DemosaicFormer] in the first part of their schedule, which is to synthetically generate rotated and flipped HybridEVS mosaics from MIPI’s ground truth images. Following BMTNet[BMTNet], we evaluate on synthetically generated quad-bayer HybridEVS mosaics, which are generated from the Kodak[Kodak], McMaster[McMaster], BSD100[BSD100] and Urban100[Urban100] datasets, as well as the first hundred images of the Waterloo Exploration Database (WED)[WED]. CFA appending is utilized for both JD3Net-S and JD3Net.
Against DemosaicFormer[DemosaicFormer], the previous state-of-the-art method, JD3Net achieves comparable performance with 3.8 times fewer FLOPs, scoring better overall on Kodak, BSD100, and WED, and scoring better SSIM on McMaster and Urban100. Averaged across all 5 datasets, JD3Net improves by 0.17 PSNR compared to DemosaicFormer while being 3.8 times faster. JD3Net-S outperforms TSANet-S by 0.08 PSNR while being 48% faster. Quantitative results are shown in Tab. 5. Some qualitative results of DemosaicFormer and JD3Net are shown in Fig. 5, which demonstrates the improved reconstruction capability of JD3Net, in particular in reconstructing the original color.
@@ 4.4 Synthetic Bayer Demosaicing
We also evaluate JD3Net on synthetic bayer demosaicing with the Kodak[Kodak] and McMaster[McMaster] datasets. Following GRL[GRL], we train both networks on the ImageNet train split. JD3Net is trained on 128x128 patches at a batch size of 20, while JD3Net-S is trained on 129x129 patches (for divisibility by 3) at a batch size of 32, both for $10^{6}$ iterations. As before, we use PSNR loss and Adam optimizer with a learning rate of $0.001$, with random flipping and rotation augmentations. CFA appending is used for JD3Net-S, but since a 4x downsampling can learn the CFA pattern inherently, CFA appending is not necessary for JD3Net.
@@
DeepJoint[DeepJoint] 14 42.00 39.14
RNAN[RNAN] 162 43.16 39.70
DRUNet[DPIR] 278 42.68 39.39
GRL-S[GRL] 491 43.57 40.22

### v2

@@ 1 Introduction
Demosaicing is a first and crucial step in an Image Sensor Processing (ISP) pipeline [7]. Recent advances in mobile photography, including more challenging non-Bayer CFA mosaics and pixel binning techniques, have made the task of demosaicing in modern mobile phones increasingly difficult [28]. Thus, deep learning techniques for non-Bayer demosaicing have been increasingly explored [23, 24, 28]. However, deep learning techniques are infamously compute-heavy, which poses challenges for usage on a mobile platform. Consequently, lightweight networks are essential for demosaicing on mobile phones [38].
One major class of image demosaicing network are isotropic networks, which are often called residual-in-residual networks [30, 28, 15]. These networks are characterized by their entire trunk maintaining the same spatial resolution; one particularly well-known example is the Vision Transformer (ViT) architecture[4]. However, unlike ViT networks, which perform aggressive patch-ification to reduce computational cost, isotropic networks for demosaicing largely perform their computations on the full spatial resolution of the image (or nearly so), consequently leading to a comparatively higher computational cost. This particular property of isotropic image demosaicing networks seems to be inherited from the strong success of isotropic networks in super-resolution[35, 13, 2]. On the contrary, for the task of image demosaicing, we find that spatial downsampling is effective for improving performance and efficiency in FLOP-equivalent network variants.
Reasoning about the effect of singular variables in designing deep networks can often be challenging, due to the dozens of interdependent hyperparameters found in network architectures. We employ two measures to enable empirical testing of the effects of downsampling. First, we consider only very simple fully-convolutional isotropic networks, which reduces our search space to only three variables: depth, width, and downsampling factor. Next, to produce principled and fair candidate variants of downsampled and non-downsampled networks, we employ a mathematical architecture design technique modified from DeepMAD[25]. Our results show that downsampling improves the performance of two FLOP-equivalent networks, especially in particularly small networks.
@@
Recent work in image demosaicing has explored complex patterns[23, 24, 38] and techniques to create unified models[11, 28]. Following this diversity in recent image demosaicing work, we test the applicability of downsampled networks to a variety of tasks. Our downsampled networks, which we term JD3Net (Joint Demosaicing and Denoising with Downsampled Network), are highly simple fully convolutional networks, but nonetheless we find that they achieve state-of-the-art efficiency on unified Bayer/non-Bayer joint-demosaicing-and-denoising, classical Bayer image demosaicing, and quad-Bayer HybridEVS image demosaicing (see Fig. 1). An overview of the CFA patterns considered can be found in Fig. 2. Overall, we argue that downsampling can broadly improve the efficiency of deep networks across a variety of image demosaicing tasks.
@@ 2.1 Tasks in Demosaicing
Joint demosaicing and denoising (JDD) is a deep learning task that combines two image processing steps – demosaicing and denoising – into one end-to-end objective. JDD is attractive because it preserves the decorrelated noise of raw images while maintaining the full information of the raw image for demosaicing[5]; consequently, research in this direction has existed even before the popularization of deep learning [9]. Nonetheless, deep learning approaches have been the mainstream in JDD research recently. Typically, a convolutional neural network (CNN) variant is applied to this task [15, 23, 24, 28, 30], with some exceptions [31, 38].
While early JDD research was largely based on the traditional Bayer CFA [30, 15, 20], modern research often focuses on non-Bayer demosaicing[23, 24, 37, 38] or unified Bayer/non-Bayer demosaicing[11, 28]. Non-Bayer CFAs tend to be more challenging, and can even have pixels with completely zero information, such as in HybridEVS image demosaicing. The difficulty of non-Bayer demosaicing, even without the denoising task, similarly merits deep learning.
@@ 2.2 Isotropic Networks for Demosaicing
The neural networks applied to demosaicing can mainly be subdivided into two types of macro-architectures: U-Net [23, 24, 31, 38, 11, 37] and Isotropic (often called residual-in-residual)[15, 30, 28, 5, 36, 12].
In general, the main difference between U-Net and isotropic macro-architectures lies in downsampling. U-Net networks employ progressive downsampling to capture multiple feature-scales [21], while the trunks of isotropic networks employ a uniform spatial size [28, 30, 35], and employ spatial downsampling or upsampling only at the beginning or end of the network.
For most general image restoration tasks, such as image denoising or super-resolution, an isotropic network employs no downsampling at all [35, 13]. For demosaicing, while many isotropic networks also do not perform downsampling [28, 12, 36], some networks perform ”packing”, which folds a CFA pattern into a single pixel[5, 30]. However, prior networks do not explicitly use downsampling to improve efficiency and do not exceed 2x downsampling. A brief summary of prior isotropic networks for demosaicing is shown in Tab. 1.
@@
DeepJoint[5] 2016 2 14
MMNet[10] 2018 1 18
RNAN [36] 2019 1 162
TENet [20] 2019 2 3199
SGNet [15] 2020 2 -
JDNDM[30] 2021 2 408
GRL[12] 2023 1 491
ESUM[28] 2025 1 408
@@
Because isotropic networks tend to operate on high spatial resolutions, without employing spatial downsampling, they are often quite slow and inefficient[2]. This is a notable problem for demosaicing, since its main application field is mobile phones, which are famously compute constrained. In particular, isotropic networks for JDD can be several hundred G-FLOPs for 256-by-256 pixel images[28, 30], which are several orders of magnitude larger than networks typically designed for image recognition on mobile devices[22, 6].
@@ 2.3 Zero-Shot Neural Architecture Search
The primary tool we will use to make this argument is Zero-Shot Neural Architecture Search (NAS). In contrast to conventional NAS techniques, Zero-Shot NAS attempts to find a performant neural network architecture without any training, often based on mathematical observations regarding the nature of deep neural networks[25, 14, 26]. This is attractive for designing networks in a principled manner, and is the primary reason we employ Zero-Shot NAS techniques in this paper.
Often, Zero-Shot NAS techniques employ the Principle of Maximum Entropy, which aims to maximize the entropy of the feature map outputted by the network [25, 14]. In particular, this paper focuses on the DeepMAD[25] entropy formulation, which is designed for convolutional neural networks and takes the following form:
@@ 3.1 Network Design
JD3Net employs modified variants of NAFBlocks[1], with the attention removed, and which we will refer to as Simplified-NAFBlocks. A specific illustration of NAFBlocks and Simplified-NAFBlocks can be found in Fig. 3.
Our decision to remove channel attention from JD3Net is due to the implementation complexity of attention in image restoration networks. For instance, techniques such as Test-time Local Converter[3] are used by some networks, but not others. Moreover, we provide some brief ablation in Sec. 4.5, which shows that the Simple Channel Attention (SCA) mechanism used by NAFNet actually decreases the performance of JD3Net. Since JD3Net can acheive strong empirical performance without complex attention mechanisms, we leave the investigation of attention for downsampled isotropic networks to future work.
@@ 3.2 Mathematical Architecture Design
As previously stated, we modify DeepMAD’s entropy score to design JD3Net. The principal issue with applying DeepMAD directly to image restoraiton is that its entropy score is reliant on the spatial resolution of the input image (ie a network would have a different entropy score with a 256x256 image compared to a 512x512 image). Consequently, DeepMAD selects different networks depending on the input resolution. For classification, where input resolution is an active consideration in network design [27, 16], this makes sense. But for image restoration, where networks should be able to process an arbitrary image, we would prefer our selection of network to be invariant to the input resolution.
@@ 4.1 Searching for JD3Net
We perform our search for JD3Net by comparing validation PSNRs on the Hard Demosaicing Dataset[28], a dataset of 638 real world joint-demosaicing-and-denoising images from 17 difficult scenes. Specifically, we compare results on the most noisy setting, ISO 3200. We train for 150 epochs using the provided pre-processed 48x48 hard training patches. Learning rate is $0.001$ on Adam optimizer with MSE loss (following ESUM[28]). Batch size for the smaller network is 64, and 48 for the larger network.
@@
JDNDM[30] 408 53.69 - - 52.34 - - 50.43 - - 49.00 - - -
BJDD[23] 69 - 50.86 - - 50.05 - - 48.88 - - 47.50 - -
SAGAN[24] 342 - - 49.55 - - 49.06 - - 48.05 - - 46.88 -
KLAP[11] 106 53.27 52.01 50.44 51.91 50.99 49.79 50.28 49.40 48.59 48.95 48.10 47.20 50.08
ESUM[28] 408 53.75 52.68 51.96 52.17 51.36 50.76 50.64 50.01 49.46 48.98 48.57 48.11 50.70
@@
BMTNet*[38] 7 37.69/0.980 34.79/0.950 36.11/0.981 34.45/0.973 33.95/0.965 35.398/0.970
NAFNet[1](w-32) 32 38.60/0.980 36.18/0.961 37.12/0.985 35.63/0.978 35.24/0.972 36.554/0.975
TSANet-S[37] 37 38.73/0.984 - 36.56/0.984 36.15/0.980 35.19/0.973 36.658/0.980
@@
Restormer[32] 282 39.16/0.986 36.54/0.967 37.11/0.985 36.36/0.977 35.00/0.971 36.83/0.977
SAGAN[24] 342 36.14/0.959 32.58/0.939 30.53/0.931 29.89/0.946 28.22/0.917 31.47/0.938
TSANet-L[37] 149 39.40/0.986 - 37.34/0.986 37.07/0.983 35.76/0.977 37.393/0.983
DemosaicFormer[31] 491 39.32/0.982 37.88/0.963 37.65/0.982 37.64/0.980 34.86/0.986 37.47/0.979
@@ 4.2 Real Image Unified Joint-Demosaicing-and-Denoising
Averaged across all CFA patterns and all noise levels, JD3Net-S achieves only 0.1 PSNR less than ESUM[28] while being 16.3 times faster, while JD3Net outperforms ESUM by 0.35 PSNR while being 3.2 times faster. We show a selection of qualitative results between JD3Net and ESUM in Fig. 4, which demonstrates the better reconstruction capability of JD3Net. Quantitative results are given in Tab. 4.
@@ 4.3 Synthetic Quad-Bayer HybridEVS Demosaicing
We also evaluate JD3Net on synthetic quad-bayer HybridEVS demosaicing. We train on the MIPI dataset [29] training split for $8\cdot 10^{6}$ iterations at a batch size of 20. We use PSNR loss[1] and Adam optimizer with a learning rate of $0.001$. We apply the data augmentations used by DemosaicFormer[31] in the first part of their schedule, which is to synthetically generate rotated and flipped HybridEVS mosaics from MIPI’s ground truth images. Following BMTNet[38], we evaluate on synthetically generated quad-bayer HybridEVS mosaics, which are generated from the Kodak[17], McMaster[34], BSD100[19] and Urban100[8] datasets, as well as the first hundred images of the Waterloo Exploration Database (WED)[18]. CFA appending is utilized for both JD3Net-S and JD3Net.
Against DemosaicFormer[31], the previous state-of-the-art method, JD3Net achieves comparable performance with 3.8 times fewer FLOPs, scoring better overall on Kodak, BSD100, and WED, and scoring better SSIM on McMaster and Urban100. Averaged across all 5 datasets, JD3Net improves by 0.17 PSNR compared to DemosaicFormer while being 3.8 times faster. JD3Net-S outperforms TSANet-S by 0.08 PSNR while being 48% faster. Quantitative results are shown in Tab. 5. Some qualitative results of DemosaicFormer and JD3Net are shown in Fig. 5, which demonstrates the improved reconstruction capability of JD3Net, in particular in reconstructing the original color.
@@ 4.4 Synthetic Bayer Demosaicing
We also evaluate JD3Net on synthetic bayer demosaicing with the Kodak[17] and McMaster[34] datasets. Following GRL[12], we train both networks on the ImageNet train split. JD3Net is trained on 128x128 patches at a batch size of 20, while JD3Net-S is trained on 129x129 patches (for divisibility by 3) at a batch size of 32, both for $10^{6}$ iterations. As before, we use PSNR loss and Adam optimizer with a learning rate of $0.001$, with random flipping and rotation augmentations. CFA appending is used for JD3Net-S, but since a 4x downsampling can learn the CFA pattern inherently, CFA appending is not necessary for JD3Net.
@@
DeepJoint[5] 14 42.00 39.14
RNAN[36] 162 43.16 39.70
DRUNet[33] 278 42.68 39.39
GRL-S[12] 491 43.57 40.22